# TrojFST: Embedding Trojans in Few-shot Prompt Tuning

### 25 Jan 2024 v2

@@ Abstract
Prompt-tuning has emerged as a highly effective approach for adapting a pre-trained language model (PLM) to handle new natural language processing tasks with limited input samples. However, the success of prompt-tuning has led to adversaries attempting backdoor attacks against this technique. Previous prompt-based backdoor attacks faced challenges when implemented through few-shot prompt-tuning, requiring either full-model fine-tuning or a large training dataset. We observe the difficulty in constructing a prompt-based backdoor using few-shot prompt-tuning, which involves freezing the PLM and tuning a soft prompt with a restricted set of input samples. This approach introduces an imbalanced poisoned dataset, making it susceptible to overfitting and lacking attention awareness. To address these challenges, we introduce TrojFST for backdoor attacks within the framework of few-shot prompt-tuning. TrojFST comprises three modules: balanced poison learning, selective token poisoning, and trojan-trigger attention. In comparison to previous prompt-based backdoor attacks, TrojFST demonstrates significant improvements, enhancing ASR $>9\%$ and CDA by $>4\%$ across various PLMs and a diverse set of downstream tasks.
@@ 1 Introduction
Prompt-tuning has become one of the most promising methods to adapting a pre-trained language model (PLM) to processing new downstream natural language processing (NLP) tasks, particularly with only few input samples (Gu et al., 2022; Zhang et al., 2022; Ma et al., 2022; Ye et al., 2022). By freezing the PLM and training with a limited set of input samples, well-optimized few-shot prompt-tuning achieves a comparable performance to full-model fine-tuning, spanning a wide spectrum of PLM sizes and NLP tasks (Gu et al., 2022; Lester et al., 2021). The success of prompt-tuning motivates adversaries to design prompt-based backdoor attacks (Xu et al., 2022; Cai et al., 2022; Du et al., 2022; Dong et al., 2023; Mei et al., 2023). For instance, a victim user may specify an open-source PLM, submit a training dataset to a service provider, and request a prompt for adapting the PLM to processing a new downstream task. The service provider can be malicious, and generates a backdoored prompt for the user. After receiving the backdoored prompt, the user may apply it to the PLM. When a trigger appears in a maliciously-prompted input sample, the PLM mis-classifies it to a predefined target class. Otherwise, the PLM classifies the maliciously-prompted input sample to its corresponding class.
@@
Unfortunately, prior prompt-based backdoors (Xu et al., 2022; Cai et al., 2022; Du et al., 2022; Dong et al., 2023; Mei et al., 2023) cannot be implemented by few-shot prompt-tuning. Prior prompt-based backdoors require either a full-model fine-tuning (Xu et al., 2022; Mei et al., 2023; Cai et al., 2022) or a large training dataset (Du et al., 2022; Dong et al., 2023). In order to achieve a high attack success rate (ASR), BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) have to modify a nontrivial number of PLM parameters, making their backdoor designs less stealthy and vulnerable to existing backdoor detection techniques (Feng et al., 2023). Although the other prompt-based backdoor designs including PPT (Du et al., 2022) and PromptAttack (Dong et al., 2023) keep the PLM clean, and tune only a small number of prompt parameters, they require hundreds of input samples to produce a backdoored prompt that can obtain a high ASR. DecodingTrust Wang et al. (2023) tests the attacks of hand-crafted engineered prompts on GPTs, not for prompt-tuning scenarios.
We find it is difficult to build a prompt-based backdoor through few-shot prompt-tuning, i.e., freezing the PLM and training a smaller set of soft prompt parameters with few (e.g., 16-shot) input samples. Naïvely training a backdoored prompt via few-shot prompt-tuning cannot achieve both a high ASR and a high clean data accuracy (CDA) at the same time for the following reasons.
  • An Imbalanced Poisoned Dataset. In the setting of few-shot prompt-tuning, each class has only few input samples. In order to enhance the ASR of the backdoored prompt, a trigger is attached to a non-trivial number of input samples belonging to the non-target classes, and the labels of these input samples are changed to the target class. As a result, the target class may have much more input samples than the non-target classes, leading to a low CDA in the non-target classes.
  • Overfitting. Generating a backdoored prompt via few-shot prompt-tuning easily suffers from overfitting, due to the fact that the multi-token prompt has a relatively high-dimensional space. Our observation reveals that when training a $20$-token backdoored prompt, the testing loss tends to be $\sim 85\%$ higher than the training loss.
  • No Attention Awareness. During the construction of a backdoored prompt via few-shot prompt-tuning, it is challenging to force the PLM to put its attention correctly on the relevant portions of the backdoor. In cases where input samples have no trigger, the PLM may allocate excessive attention to the backdoored prompt, leading to a low CDA. Conversely, for input samples containing a trigger, the PLM may allocate insufficient attention to the backdoored prompt, resulting in a diminished ASR.
@@
In this paper, we propose a prompt-based backdoor attack, TrojFST, against PLMs through few-shot prompt-tuning. Instead of a full-model fine-tuning, TrojFST freezes the PLM, and trains a backdoored prompt for the PLM with only few input samples by tuning only one prompt token. The PLM remains untainted throughout the entirety of our TrojFST attack, making TrojFST more stealthy and resistant to existing encoder backdoor detection techniques (Feng et al., 2023). Compared to prior prompt-based backdoor attacks, TrojFST improves the ASR by $9\%\sim 48\%$ and the CDA by $4\%\sim 9\%$ across various PLMs and a wide range of downstream tasks.
@@ 2.1 Prompt-tuning for PLMs
Table 1: The comparison between TrojFST and prior prompt-based backdoors including BToP (Xu et al., 2022), Notable (Mei et al., 2023), BadPrompt (Cai et al., 2022), PPT (Du et al., 2022), and PromptAttack (Dong et al., 2023).
@@
TrojFST
2.2 The Limitations of Prior Prompt-based Backdoors
Although the success of prompt-tuning motivates adversaries to build prompt-based backdoor attacks, prior prompt-based backdoor attacks require high training costs, i.e., either expensive full-model fine-tuning (Xu et al., 2022; Mei et al., 2023; Cai et al., 2022) or a large training dataset (Du et al., 2022; Dong et al., 2023). No backdoors on prompt-tuning with frozen PLMs on few-shot downstream samples. We compare prior prompt tuning based backdoor attacks and TrojFST in Table 1. Compared to prior prompt-based backdoors, TrojFST is the only prompt-based backdoor attack implemented by few-shot prompt-tuning. Among prior prompt-based backdoors, BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) have to invoke a full-model fine-tuning on their PLMs, making themselves less stealthy and vulnerable to existing encoder backdoor detection techniques (Feng et al., 2023). Notably, although BadPrompt (Cai et al., 2022) aims to produce task-specific poisoned prompts by few input samples, it has to modify not only the continuous prompt parameters but also the PLM parameters (see Equation 1 in (Cai et al., 2022)) during its backdoor generation. Although PPT (Du et al., 2022), and PromptAttack (Dong et al., 2023) freeze the PLMs and tune only a small set of prompt parameters, they require a large training dataset consisting of hundreds of input samples. In contrast, TrojFST can generate a backdoored prompt that can achieve both a high ASR and a high CDA by freezing the PLMs and tuning a small set of prompt parameters with few (e.g., 16-shot) input samples.
2.3 The Difficulties in Building Prompt-based Backdoors via Few-shot Prompt-tuning
@@
$\vspace{0.1in}\begin{split}\mathcal{L}&=\overbrace{\sum_{(x_{i},y_{i})\in\mathcal{D}_{b}}\mathcal{L}(f(x_{i}),y_{i})}^{CDA}+\overbrace{\lambda\cdot\sum_{(x_{i},y_{i})\in\mathcal{D}_{p},i\neq t}\mathcal{L}(f(x_{i}+\tau),y_{t})}^{ASR}\\ &=\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}),y_{i})}^{\textbf{non-target class}}+\overbrace{\sum_{j=0}^{m}\mathcal{L}(f(x_{t}^{j}),y_{t})}^{\textbf{target class}}+\overbrace{\lambda\cdot\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}+\tau),y_{t})}^{\textbf{$\lambda m\alpha\cdot(n-1)$ target class samples}}\end{split}\vspace{-0.3in}$ (1)
@@
where $\mathcal{L}$ is the cross-entropy loss function, $x_{i}$ is an input sample belonging to the $i_{th}$ class, $y_{i}$ is the label of the $i_{th}$ class, $y_{t}$ represents the label of the target class, $f()$ indicates the output of the prompted PLM, $\tau$ represents a trigger, and ($x_{i}+\tau$,$y_{t}$) is a poisoned input sample. Note that a syntactic trigger does not possess an individual $\tau$, yet for the sake of general expression, we universally denote $x_{i}+\tau$ as a poisoned sample. $\mathcal{D}_{b}$ means a benign dataset, $\mathcal{D}_{p}$ indicates a poisoned dataset, and $\lambda$ denotes the weight of the ASR loss. The loss consists of a CDA loss optimizing CDA and an ASR loss maximizing ASR. The CDA loss can be further decomposed into the CDA loss for the non-target classes and the CDA loss for the target class, as shown in the second line of Equation 1, where $n$ is the class number, $m$ is the shot number, $x_{i}^{j}$ indicates the $j_{th}$ input sample belonging to the $i_{th}$ class, and $\alpha\in[0,1]$ is the percentage of the poisoned input samples in the input samples. In Equation 1, the CDA loss for the target class requires $m$ input samples to train the normal behavior of the prompt, while the ASR loss for the target class needs $\lambda m\alpha\cdot(n-1)$ input samples to train the malicious behavior of the prompt. In total, the target class receives $[m+\lambda m\alpha\cdot(n-1)]$ input samples, which is more than the $m$ input samples used to train the other non-target classes. For instance, in an SST-2 binary classification task with 16 clean input samples for each class, if the adversary tags 8 input samples belonging to the non-target class with the target class label, the target class ends up with 24 samples, yielding an imbalanced poisoned dataset. The details of our experimental methodology can be viewed in Section 4. Consequently, both the CDA of the non-target class and the overall CDA greatly decrease with an increasing number of the poisoned input samples, although the CDA of the target class and the ASR increase with more poisoned input samples.
@@
Overfitting. Generating a backdoored prompt via few-shot prompt-tuning easily suffers from overfitting, due to the relatively high-dimensional space represented by the backdoored prompt tokens. When training a $20$-token backdoored prompt to attack RoBERTa-Large on SST-2, the testing loss is $50\%\sim 85\%$ larger than the training loss.
@@
No Attention Awareness. We naïvely used few-shot prompt-tuning to build a backdoored prompt. The attention score received by the backdoored prompt when processing a clean input is very similar to that of the backdoored prompt when processing a poisoned input sample containing a trigger, indicating that the backdoored prompt generated by naïve few-shot prompt-tuning has no attention awareness. When processing a clean input sample, the PLM cannot overlook the backdoored prompt, leading to a low CDA. Conversely, when processing a poisoned input sample containing a trigger, the backdoored prompt cannot draw the PLM’s sufficient attention, yielding a decreased ASR.
@@
As Table 1 highlights, our TrojFST balances the poisoned dataset by dynamically reducing the number of input samples belonging to the predefined target class based on the number of poisoned input samples from the non-target classes. Moreover, TrojFST tunes only one token in the backdoored prompt to overcome the overfitting problem. Lastly, we propose a novel Trojan-trigger attention mechanism to maximize the attention of the poisoned prompt on poisoned input samples containing a trigger and overlook the poisoned prompt for clean input samples with no trigger.
3 TrojFST
@@ 3.1 Threat Model
Attacker’s objective. We assume an attacker can train a prompt to adapt a PLM (e.g., Google T5) to processing a downstream task by few-shot input samples and inject a backdoor into the prompt that can be activated by an invisible syntactic trigger (Qi et al., 2021). Then, a victim user receives the backdoored prompt. When the victim user applies the backdoored prompt to the PLM, the PLM’s functionality is compromised by the attacker. More specifically, the PLM acts normally with benign input samples. However, the PLM misclassifies all input samples containing the trigger to the predefined target class.
@@
3.2 Balancing the Poisoned Dataset
In the setting of few-shot prompt-tuning, every class initially has an equal number of clean input samples. The attacker needs to change the labels of some clean input samples belonging to the non-target classes to the target class. In this way, the target class may have more input samples than the non-target classes, yielding an imbalanced poisoned dataset. The accuracy of the non-target class and the overall accuracy greatly decrease as more poisoned input samples are inserted into the target class.
To mitigate the imbalanced poisoned dataset issue, one possible solution is to decrease the value of $\lambda$ in Equation 1. However, when $\lambda$ is not zero, the target class sample number is still larger than the non-target class, and a smaller $\lambda$ yields only a lower ASR. For instance, when setting $\lambda=0.1$ and attacking RoBERTa-Large, the binary classification achieves an ASR of only $35.39\%$ on SST-2. Therefore, decreasing the value of $\lambda$ cannot solve the problem of the imbalanced poisoned dataset.
We propose a balanced poison learning technique to reduce the number of clean input samples in the target class (i.e., $m$) during the process of data poisoning. We add a corrective factor denoted as $\beta$ ($\beta\in(0,1)$) to the CDA loss item of the clean samples belonging to the target class. Our new backdoor loss can be summarized as follows:
$\mathcal{L}=\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}),y_{i})}^{\textbf{non-target class}}+\overbrace{\mathbf{\beta}\cdot\sum_{j=0}^{m}\mathcal{L}(f(x_{t}^{j}),y_{t})}^{\textbf{target class}}+\overbrace{\lambda\cdot\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}+\tau),y_{t})}^{\textbf{$\lambda m\alpha\cdot(n-1)$ target class samples}}$ (2)
This modification on the backdoor loss ensures that the number of input samples belonging to the target class is equal to that belonging to each non-target class, i.e., $m\cdot\beta+\lambda m\alpha\cdot(n-1)=m\Rightarrow\beta+\lambda\alpha\cdot(n-1)=1$. For a given set of $\lambda$ and $\alpha$ configurations, we can adjust $\beta$ to maintain the equality. We studied the impact of various configurations of $\lambda$, $\alpha$, and $\beta$ for TrojFST in Table 5.15.1.

### 19 Mar 2024 v3

# TrojFSP: Trojan Insertion in Few-shot Prompt Tuning

@@ Abstract
Prompt tuning is one of the most effective solutions to adapting a fixed pre-trained language model (PLM) for various downstream tasks, especially with only a few input samples. However, the security issues, e.g., Trojan attacks, of prompt tuning on a few data samples are not well-studied. Transferring established data poisoning attacks directly to few-shot prompt tuning presents multiple challenges. One significant issue is the poisoned imbalance issue, where non-target class samples are added to the target class, resulting in a greater number of target-class samples compared to non-target class. While this issue is not critical in regular tuning, it significantly hampers the few-shot prompt tuning, making it difficult to simultaneously achieve a high attack success rate (ASR) and maintain clean data accuracy (CDA). Additionally, few-shot prompting is prone to overfitting in terms of both ASR and CDA. In this paper, we introduce TrojFSP, a method designed to address the challenges. To solve the poisoned imbalance issue, we develop a Target-Class Shrink (TC-Shrink) technique, which aims to equalize the number of poisoning samples. To combat overfitting, we employ a Selective Token Poisoning technique to boost attack performance. Furthermore, we introduce a Trojan-Trigger Attention objective function to amplify the attention of the poisoned trojan prompt on triggers. Experiments show that our TrojFSP achieves an ASR of over 99% while maintaining negligible decreases in CDA across various PLMs and datasets.
TrojFSP: Trojan Insertion in Few-shot Prompt Tuning
Mengxin Zheng1thanks: Corresponding Author Email: mengxin.zheng@ucf.edu. ,      Jiaqi Xue1,       Xun Chen2 YanShan Wang3,      Qian Lou1,       Lei Jiang4 1University of Central Florida 2Samsung Research America 3University of Pittsburgh 4Indiana University Bloomington
@@ 1 Introduction
Prompt-tuning has become one of the most promising methods to adapting a pre-trained language model (PLM) to processing new downstream natural language processing (NLP) tasks, particularly with only few input samples (Gu et al., 2022; Zhang et al., 2022; Ma et al., 2022; Ye et al., 2022). By freezing the PLM and training with a limited set of input samples, well-optimized few-shot prompt-tuning achieves a comparable performance to full-model fine-tuning, spanning a wide spectrum of PLM sizes and NLP tasks (Gu et al., 2022; Lester et al., 2021). The success of prompt-tuning motivates adversaries to design prompt-based Trojan (a.k.a backdoor) attacks (Xu et al., 2022; Cai et al., 2022; Du et al., 2022; Shi et al., 2022; Mei et al., 2023; Xue et al., 2024). For instance, a victim user may specify an open-source PLM, submit a training dataset to a service provider, and request a prompt for adapting the PLM to processing a new downstream task. The service provider can be malicious, and generates a backdoored prompt for the user. After receiving the backdoored prompt, the user may apply it to the PLM. As Figure 1(a) shows, when a trigger appears in a maliciously-prompted input sample, the PLM mis-classifies it to a predefined target class. Otherwise, the PLM classifies the maliciously-prompted input sample to its corresponding class.
Figure 1: The overview of our proposed TrojFSP attack.
@@
Unfortunately, prior prompt-based backdoors (Xu et al., 2022; Cai et al., 2022; Du et al., 2022; Shi et al., 2022; Mei et al., 2023) cannot be implemented by few-shot prompt-tuning. Prior prompt-based backdoors require either a full-model fine-tuning (Xu et al., 2022; Mei et al., 2023; Cai et al., 2022) or a large training dataset (Du et al., 2022; Shi et al., 2022). In order to achieve a high attack success rate (ASR), BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) have to modify a nontrivial number of PLM parameters, making their backdoor designs less stealthy and vulnerable to existing backdoor detection techniques (Feng et al., 2023; Zheng et al., 2023b). Although the other prompt-based backdoor designs including PPT (Du et al., 2022) and PromptAttack (Shi et al., 2022) keep the PLM clean, and tune only a small number of prompt parameters, they require hundreds of input samples to produce a backdoored prompt that can obtain a high ASR. DecodingTrust Wang et al. (2023) evaluates the effectiveness of attacks using hand-crafted, engineered prompts on GPT models. However, it does not address scenarios involving prompt-tuning.
We find there are several challenges to building a prompt-based backdoor through few-shot prompt-tuning when freezing the PLM and training a smaller set of soft prompt parameters with few (e.g., 16-shot) input samples. Naïvely training a backdoored prompt via few-shot prompt-tuning cannot achieve both a high ASR and high clean data accuracy (CDA) at the same time for the following reasons. (i) poisoned imbalance issue. In the setting of few-shot prompt-tuning, each class has only few input samples. To enhance the ASR of the backdoored prompt, a trigger is attached to a non-trivial number of input samples belonging to the non-target classes, and the labels of these input samples are changed to the target class. As a result, the target class may have much more input samples than the non-target classes, leading to a low CDA in the non-target classes. (ii) ASR and CDA Overfitting. Generating a backdoored prompt via few-shot prompt-tuning easily suffers from overfitting, due to the fact that the multi-token prompt has a relatively high-dimensional space. Our observation reveals that when training a $20$-token backdoored prompt, the testing loss tends to be $\sim 85\%$ higher than the training loss. (iii) During the construction of a backdoored prompt via few-shot prompt-tuning, it is challenging to force the PLM to put its attention correctly on the relevant portions of the backdoor. In cases where input samples have no trigger, the PLM may allocate excessive attention to the backdoored prompt, leading to a low CDA. Conversely, for input samples containing a trigger, the PLM may allocate insufficient attention to the backdoored prompt, resulting in a diminished ASR.
@@
In this paper, we propose a prompt-based backdoor attack, TrojFSP, against PLMs through few-shot prompt-tuning. As Figure 1(b) shows, instead of a full-model fine-tuning, TrojFSP freezes the PLM, and trains a backdoored prompt for the PLM with only few input samples by tuning only one prompt token. The PLM remains untainted throughout the entirety of our TrojFSP attack, making TrojFSP more stealthy and resistant to existing encoder backdoor detection techniques (Feng et al., 2023). Compared to prior prompt-based backdoor attacks, TrojFSP improves the ASR by $9\%\sim 48\%$ and the CDA by $4\%\sim 9\%$ across various PLMs and a wide range of downstream tasks.
@@ 2.1 Prompt-tuning for PLMs
Table 1: The comparison between TrojFSP and prior prompt-based backdoors including BToP (Xu et al., 2022), Notable (Mei et al., 2023), BadPrompt (Cai et al., 2022), PPT (Du et al., 2022), and PromptAttack (Shi et al., 2022).
@@
TrojFSP
2.2 The Limitations of Prior Attacks
Backdoor attack is one of the most dangerous malicious attacks Gu et al. (2017); Xue and Lou (2022); Jia et al. (2022); Zheng et al. (2023a); Xue et al. (2024). In a backdoor attack, a backdoor is injected into a neural network model, allowing the model to behave normally for benign inputs but inducing a predefined behavior for any inputs with a trigger. With the success of prompt-tuning, adversaries design prompt-based backdoor attacks. However, prior prompt-based backdoor attacks require high training costs, i.e., either expensive full-model fine-tuning (Xu et al., 2022; Mei et al., 2023; Cai et al., 2022) or a large training dataset (Du et al., 2022; Shi et al., 2022). We compare prior prompt tuning based backdoor attacks and TrojFSP in Table 1. Compared to prior prompt-based backdoors, TrojFSP is the only prompt-based backdoor attack implemented by few-shot prompt-tuning. Among prior prompt-based backdoors, BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) have to invoke a full-model fine-tuning on their PLMs, making themselves less stealthy and vulnerable to existing encoder backdoor detection techniques (Feng et al., 2023). Notably, although BadPrompt (Cai et al., 2022) aims to produce task-specific poisoned prompts by few input samples, it has to modify not only the continuous prompt parameters but also the PLM parameters (see Equation 1 in (Cai et al., 2022)) during its backdoor generation. Although PPT (Du et al., 2022), and PromptAttack (Shi et al., 2022) freeze the PLMs and tune only a small set of prompt parameters, they require a large training dataset consisting of hundreds of input samples. In contrast, TrojFSP can generate a backdoored prompt that can achieve both a high ASR and a high CDA by freezing the PLMs and tuning a small set of prompt parameters with few (e.g., 16-shot) input samples.
2.3 Motivation
@@
$\begin{split}\mathcal{L}&=\overbrace{\sum_{(x_{i},y_{i})\in\mathcal{D}_{b}}\mathcal{L}(f(x_{i}),y_{i})}^{CDA}+\overbrace{\sum_{(x_{i},y_{i})\in\mathcal{D}_{p},i\neq t}\mathcal{L}(f(x_{i}+\tau),y_{t})}^{ASR}\\ &=\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}),y_{i})}^{\textbf{non-target class}}+\overbrace{\sum_{j=0}^{m}\mathcal{L}(f(x_{t}^{j}),y_{t})}^{\textbf{target class}}\\ &\quad+\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}+\tau),y_{t})}^{\textbf{$m\alpha\cdot(n-1)$ target class samples}},\end{split}$ (1)
@@
where $\mathcal{L}$ is the cross-entropy loss function, $x_{i}$ is an input sample belonging to the $i_{th}$ class, $y_{i}$ is the label of the $i_{th}$ class, $y_{t}$ represents the label of the target class, $f()$ indicates the output of the prompted PLM, $\tau$ represents a trigger, and ($x_{i}+\tau$, $y_{t}$) is a poisoned input sample. Note that a syntactic trigger does not possess an individual $\tau$, yet for the sake of general expression, we universally denote $x_{i}+\tau$ as a poisoned sample. $\mathcal{D}_{b}$ means a benign dataset, $\mathcal{D}_{p}$ indicates a poisoned dataset. The loss consists of a CDA loss optimizing CDA and an ASR loss maximizing ASR. The CDA loss can be further decomposed into the CDA loss for the non-target classes and the CDA loss for the target class, as shown in the second line of Equation 1, where $n$ is the class number, $m$ is the shot number, $x_{i}^{j}$ indicates the $j_{th}$ input sample belonging to the $i_{th}$ class, and $\alpha\in[0,1]$ is the percentage of the poisoned input samples in the input samples. In Equation 1, the CDA loss for the target class requires $m$ input samples to train the normal behavior of the prompt, while the ASR loss for the target class needs $m\alpha\cdot(n-1)$ input samples to train the malicious behavior of the prompt. In total, the target class receives $[m+m\alpha\cdot(n-1)]$ input samples, which is more than the $m$ input samples used to train the other non-target classes. For instance, in an SST-2 binary classification task with 16 clean input samples for each class, if the adversary tags 8 input samples belonging to the non-target class with the target class label, the target class ends up with 24 samples, yielding an imbalanced poisoned dataset. The details of our experimental methodology can be viewed in Section 4. Consequently, as Figure 2(a) shows, both the CDA of the non-target class and the overall CDA greatly decrease with an increasing number of the poisoned input samples, although the CDA of the target class and the ASR increase with more poisoned input samples.
@@
Overfitting. Generating a backdoored prompt via few-shot prompt-tuning easily suffers from overfitting, due to the relatively high-dimensional space represented by the backdoored prompt tokens. As Figure 2(b) shows, when training a $20$-token backdoored prompt to attack RoBERTa-Large on SST-2, the testing loss is $50\%\sim 85\%$ larger than the training loss.
@@
No Attention Awareness. We naïvely used few-shot prompt-tuning to build a backdoored prompt. As Figure 2(c) shows, the attention score received by the backdoored prompt when processing a clean input is very similar to that of the backdoored prompt when processing a poisoned input sample containing a trigger, indicating that the backdoored prompt generated by naïve few-shot prompt-tuning has no attention awareness. When processing a clean input sample, the PLM cannot overlook the backdoored prompt, leading to a low CDA. Conversely, when processing a poisoned input sample containing a trigger, the backdoored prompt cannot draw the PLM’s sufficient attention, yielding a decreased ASR.
@@
As Table 1 highlights, our TrojFSP balances the poisoned dataset by dynamically reducing the number of input samples belonging to the predefined target class based on the number of poisoned input samples from the non-target classes. Moreover, TrojFSP tunes only one token in the backdoored prompt to overcome the overfitting problem. Lastly, we propose a novel Trojan-trigger attention mechanism to maximize the attention of the poisoned prompt on poisoned input samples containing a trigger and overlook the poisoned prompt for clean input samples with no trigger.
Figure 2: The difficulties in the backdoor attack via few-shot prompt-tuning (SST-2 with RoBERTa-Large model): (a) an imbalance poisoned dataset. (b) overfitting. (c) no attention awareness.
3 TrojFSP
@@ 3.1 Threat Model
Attacker’s objective. We assume an attacker can train a prompt to adapt a PLM (e.g., Google T5) to processing a downstream task by few-shot input samples and inject a backdoor into the prompt that can be activated by an invisible syntactic trigger (Qi et al., 2021; Lou et al., 2022; Al Ghanim et al., 2023). Then, a victim user receives the backdoored prompt. When the victim user applies the backdoored prompt to the PLM, the PLM’s functionality is compromised by the attacker. More specifically, the PLM acts normally with benign input samples. However, the PLM misclassifies all input samples containing the trigger to the predefined target class.
@@
3.2 Target-Class Shrink (TC-Shrink)
In the setting of few-shot prompt-tuning, every class initially has an equal number of clean input samples. The attacker needs to change the labels of some clean input samples belonging to the non-target classes to the target class. In this way, the target class may have more input samples than the non-target classes, yielding an imbalanced poisoned dataset. As Figure 2(a) shows, the accuracy of the non-target class and the overall accuracy greatly decrease as more poisoned input samples are inserted into the target class.
To mitigate the imbalanced poisoned dataset issue, one possible solution is to decrease the value of $\alpha$ in Equation 1. However, when $\alpha$ is not zero, the target class sample number is still larger than the non-target class, and a smaller $\alpha$ yields only a lower ASR. For instance, when setting $\alpha=0.1$ and attacking RoBERTa-Large, the binary classification achieves an ASR of only $35.39\%$ on SST-2. Therefore, decreasing the value of $\alpha$ cannot solve the problem of the imbalanced poisoned dataset.
We propose Target-Class Shrink (TC-Shrink) technique to reduce the number of clean input samples in the target class (i.e., $m$) during the process of data poisoning. We add a corrective factor denoted as $\beta$ ($\beta\in(0,1)$) to the CDA loss item of the clean samples belonging to the target class. Our new backdoor loss can be summarized as follows:
$\begin{split}\mathcal{L}&=\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}),y_{i})}^{\textbf{non-target class}}+\overbrace{\mathbf{\beta}\cdot\sum_{j=0}^{m}\mathcal{L}(f(x_{t}^{j}),y_{t})}^{\textbf{target class}}\\ &+\overbrace{\sum_{i\neq t}^{n}\sum_{j=0}^{m}\mathcal{L}(f(x_{i}^{j}+\tau),y_{t})}^{\textbf{$m\alpha\cdot(n-1)$ target class samples}}\end{split}$ (2)
This modification on the backdoor loss ensures that the number of input samples belonging to the target class is equal to that belonging to each non-target class, i.e., $m\cdot\beta+m\alpha\cdot(n-1)=m\Rightarrow\beta+\alpha\cdot(n-1)=1$. For a given set of $\beta$ configurations, we can adjust $\alpha$ to maintain the equality. We studied the impact of various configurations of $\alpha$, and $\beta$ for TrojFSP in Table 5.1.